a) Given the following simple Keynesian Model: Y = C + I + G + X-M, 

where 

Consumption schedule is given as C= 100 +0.75Y

Investment (I) = 50

Government (G) = 100

and Net Export (X-M) = 20

i. Calculate the Equlibrium Level of Income [4 Marks]

ii. Calculate the size of Consumption at the Equilibrium Level



1
Expert's answer
2021-05-18T12:51:22-0400

i)

Using the equilibrium condition in the Keynesian Model:

"Y = C+ I+ G + NX"substituting the given values

"Y=(100 +0.75Y) + 50 +100 + 20," simplifying it:

"Y=270+0.75Y,"

"Y-0.75=270"

"0.25Y=270"

"Y=\frac{270}{0.25}=1080"

Thus, the equilibrium level of income amounts to 1080


ii)

At Y equals to 1080, the size of consumption is:


"C= 100 +0.75Y, put Y =1080"

"=100 +0.75\u00d71080"

"=100+810"

"=910"

Hence, the consumption expenditure amounts to 910.